HYPOTHESIS TESTINGABOUT POPULATION MEAN š

USING THE CRITICAL VALUE APPROACH


The records of SCA Registrar show that the average final grade in Mathematics for STEM students is 91 with a standard deviation of 3. A group of student-researchers found out that the average final grade of 37 randomly selected STEM students in Mathematics is no longer 91. Use 0.05 level of significance to test the hypothesis and a sample mean within the range of 88 to 94 only.

A. State the hypotheses.

B. Determine the test statistic to use.

C. Determine the level of significance, critical value, and the decision rule.

D. Compute the value of the test statistic.

E. Make a decision.

F. Draw a conclusion.



1
Expert's answer
2022-05-31T08:40:46-0400

A. The following null and alternative hypotheses need to be tested:

"H_0:\mu=91"

"H_1:\mu\not=91"


B.This corresponds to a two-tailed test, for which a z-test for one mean, with known population standard deviation will be used.


C. Based on the information provided, the significance level isĀ "\alpha = 0.05,"Ā and the critical value for a two-tailed test isĀ "z_c = 1.96."

The rejection region for this two-tailed test isĀ "R = \{z:|z|>1.96\}."


D. The z-statistic is computed as follows:



"z=\dfrac{\bar{x}-\mu}{\sigma/\sqrt{n}}=\dfrac{90-91}{3/\sqrt{37}}=-2.0276"


E. Since it is observed thatĀ "|z|=2.0276>1.96=z_c,"Ā it is then concluded thatĀ the null hypothesis is rejected.

Using the P-value approach:

The p-value isĀ "p=2P(z<-2.0276)= 0.042601,"Ā and sinceĀ "p= 0.042601<0.05=\alpha,"Ā it is concluded that the null hypothesis is rejected.


F. Therefore, there is enough evidence to claim that the population meanĀ "\mu"

is different than 91, at theĀ "\alpha = 0.05"Ā significance level.